OpenStudy (anonymous):

x-y=3 2x-2y=6 substitution

OpenStudy (anonymous):

x-y=3(Multiply with - 2) 2x-2y=6 -2x+2y=-6 2x-2y=6 ======== They all will be cancelled so no solution.

OpenStudy (anonymous):